You purchased a home with a mortgage of $275,000. Payments on the mortgage are $1,400 at the beginning of each month and the annual interest rate on the mortgage is 5.50%. How many years will it take to pay off your mortgage? Round your answer to 2 decimals.
Answer #1

Solution: Calculation of number of years required to repay the mortgage.

Following are the steps to be followed on Microsoft Excel to calculate the Number of years:

Step 1: Click on "FORMULAS" tab at the top of Microsoft Excel
Step 2: Select the option "Financial"
Step 3: Under "Financial" select the option "NPER"
Step 4: Insert Rate = 0.055/12    PMT = -1400 PV = 275000   TYPE = 1

NPER = 495.356

Number of months required to pay off the mortgage is 495.356 Months

Number of years required to pay off the mortgage = 495.356/12 = 41.28 Years